Operations Report – February 9th

Crew 188 Operational Report 09/02/2018

Sol 12

Zac Trolley

Non-Nominal Systems: electrical system, electrical cords, propane, EVA Suits

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 14.5h

Solar: SOC 80%

Diesel: 72%

Propane: 55

Ethanol Free Gasoline (5 Gallon containers for ATV): 3 Gallons

Water (trailer): 0 Gallons

Water (static): 65 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Meter: 130949.8

Toilet tank emptied: No

ATVs Used: Yes

Oil Added: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0.5

Notes on ATVs: The Gasoline storage area needs to be moved. The sun is creating a large pressure differential on the jerry cans, making that area an explosion risk.

Deimos rover used: No

Hours: Hours: 108.1

Beginning charge:

Ending charge:

Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 19.5h

Beginning charge:

Ending charge:

Currently charging: No

Opportunity rover used: No

Hours: 19.1h

Beginning charge:

Ending charge:

Currently charging: No

Curiosity rover used: No

Hours: 12.0h

Beginning charge:

Ending charge:

Currently charging: No

HabCar: No

General notes and comments:

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Generator was turned off at 9:25

Generator was turned on at 18:24

I’ve found two extension cords that have been damaged and need to be replaced. If there are 3 prong plugs available, I can repair the cords. Please advise.

Summary of internet: Tests were run on the internet connectivity and passed on to Bernard.

Summary of suits and radios: I’ve passed on a report about the EVA suits

Summary of Hab operations: Propane was smelt again this afternoon around 4pm. The heater had been turned off for 24 hours, and nothing else was on. I strongly suggest a licensed contractor investigate the propane leak before another team moves into Hab. There is a significant safety risk as we have not been able to pinpoint the source of the leak.

Summary of EVA operations: Nominal

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: We could use another bag of toilet deodorizer, as we only have one puck left.

Operations Report – February 4th

Crew 188 Operational Report 04/02/2018

Sol 7

Zac Trolley

Non-Nominal Systems: electrical system, electrical cords

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 16.9h

Solar: SOC 80%

Diesel: 75%

Propane: 61 psi

Ethanol Free Gasoline (5 Gallon containers for ATV): 8 Gallons

Water (trailer): 0 Gallons

Water (static): 170 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Meter 130739.8

Toilet tank emptied: No

ATVs Used: None

Oil Added: None

ATV Fuel Used: None

# Hours the ATVs were Used today: 0

Notes on ATVs: none

Deimos rover used: No

Hours: Hours: 108.1

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 18.0h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: No

Hours: 13.5

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Curiosity rover used: No

Hours: 11.3h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

HabCar: Not used

General notes and comments: We are rotating the Rovers with the one plug in we have, and that should be sufficient if we are taking 2 rovers at a time per EVA.

It would seem that the electrical generation isn’t being regulated that well. The diesel and solar power systems will vary between 120V and 140V. No known reason why this is happening.

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Generator was turned off at 10:36

Generator was turned on at 17:28

I’ve found two extension cords that have been damaged and need to be replaced. If there are 3 prong plugs available, I can repair the cords. Please advise.

Summary of internet: Internet connectivity continues to be a problem even with the data purchased. Devices are having troubles connecting and this is limiting the crew’s ability to monitor the CapCom emails

Summary of suits and radios: The clasp on Suit 1 was bent back to place, it’s as good as new.

Suit 2 was repaired, it was a cold solder joint, as well as the connection not being crimped. I expertly re-attached the connector and the suit is operational

Summary of Hab operations: Nominal

Summary of EVA operations: In the warm weather the fogging issue isn’t as pronounced. We are hoping for warm weather to maintain visual accuracy in the helmets.

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: We could use another bag of toilet deodorizer, as we only have one pack left.

EVA Report – February 9th

Crew 188 EVA #14 Summary Report 09FEB2018

Author’s name: Julia DeMarines, Crew Astronomer, Green Hab Operator

Purpose of EVA: To perform additional video filming using either the GoPRO or Insta360 PRO camera on the ridge around the Hab. While walking, we will collect rock and plant samples for in-situ chlorophyll samples for analysis.

Location of EVA: Hab Ridge
UDM27 Coordinates: 12S, 5175750E, 54250500N

Number of EVA Crew: 3

Participants: Ryan Kobrick, Sarah Jane Pell, Julia DeMarines

EVA Commander: Julia DeMarines




Road(s) and routes per MDRS map:


1) Exit Hab and head to the hills directly west to the Hab

2) Walk to the Sagan Road


3) Climb to summit Hab Ridge

4) Follow the road on Hab Ridge Road as time allows

Mode of Travel: Walking

Vehicles you will be using (if applicable): None.

Duration: 2 hours 41 minutes


EVA Departure Time: 10:26


EVA Return Time: 13:07

Summary:

We left the hab a little after the predicted time due to two last minute bathroom needs. The weather was sunny with some haze and no wind. The crew soon felt overheated, helmets/glasses were fogging up and had to stop frequently to let the sun burn off the fog. Once on the ridge, Julia selected two sites for chlorophyll detection samples, and Sarah Jane set up her Insta360º Pro camera for three different shots. Ryan recorded the journey on a GoPro camera to document the experience. Once our objectives were accomplished we turned around and walked back to the hab.

Journalist Report – February 9th

MDRS Crew 188 Journalist Report 09FEB2018

“Every Sol, our soul expands”.

Sol 12: Authors’ Name: Dr. Sarah Jane Pell, Artist-in-Residence

We Crew 188 have all participated in EVA research experiments and scouting exercises, initiated in workshop and research activities, partaken in movement and reflective pursuits, cultural and culinary exchange, and interviews and candid conversations. Unintended outcomes informed an adaptation of Maslow’s human needs for future life on Mars, numerous operational recommendations and feedback documents, a range of education and public outreach activities, and the anticipated science, arts, humanities and engineering data collection for localized and on-going research. We have also been actively engaged in social media for the purpose of learning, discovering, sharing and promoting interdisciplinary exploration and Earth analogues to contribute a critical cultural and aesthetic suite of responses to the MDRS experience, but have we learned what it takes?

As the night falls, and we watch the ISS above us in the night sky, we connect our journey to the six astronauts currently in space, to the lives of those who have so generously and courageously expanded our capacity for imagination and daring through selfless research and discovery. While we are in no position to compare, we remain humbled, and inspired, in our pursuit of our dreams to be among those who contribute to space exploration. We took time out from our professional lives to come to MDRS to live day-to-day life as early Mars settlers. We navigated the terrain of the desert analogue environment and the simulation conditions. We worked as a crew. We discovered each other and reflected on aspects of our selves and our suitability for the tasks through our engagement with the daily operational and environmental challenges. More importantly, we cared for each other. We cared for the habitat and the landscape. We cared for those who came here before us, and those who are on their way: the future analogue crews and support teams. We remembered that in coming together for peaceful and united purpose, that we can accomplish small things in great ways, and great things in seemingly insignificant ways. We supplemented each other’s weaknesses and complemented each other’s strengths, and gained new perspectives through our transfer and exchange of values and experiences. We discovered what systems and approaches worked effortlessly on Mars and what took immense energy, insight, courage, patience and tact to overcome, or live with. We imagined, created, dreamed and put into action that which inspired our heartfelt curiosities and professional interests. As our cheesy Crew 188 fridge affirmation said, “Every Sol, our soul expands”.

I want to wish the in-coming Crew 189 every success and take this opportunity to personally thank the amazing people who make up Crew 188: Dr. Ryan Kobrick (KOB1), Renee Garifi (Llama Llama Ding Dong), Zac Trolley (Boltz/Bootz), Julia De Marines (Jules Verne), and Tatsunari Tomiyama (Tom Cat). I think the world of you all, and could not have imagined a better crew to explore the red plains with. Thanks also to the Mars Society Operations and Mission Support for their on-going support of our mission. A part of MDRS will be with us always.

Signing off,

Dr. Sarah Jane Pell (Bubbles/SJ), Journalist in Residence, Crew 188.

– – – – –

Dr. Sarah Jane Pell

TED Fellow 2010, Australia Council Fellow 2016, Gifted Citizen 2016

The ‘Performing Astronautics: following the Body’s Natural Edge into the Abyss of Space’ project is assisted by the Australian Government through the Australia Council, its arts funding and advisory body http://www.artistastronaut.com

Artist, Occupational Diver, Keynote Speaker, Researcher, Simulation Astronaut http://www.sarahjanepell.com

EVA Report – February 9th

EVA #15 EVA Report 09Feb2018

Author’s name: Tatsunari Tomiyama AHFP, Health & Safety Officer

Purpose of EVA: To collect micrometeorite soil samples at the Tooth Ridge and take photo and video for the mission documentary.

Location of EVA: Toothy Ridge

UDM27 Coordinates: 12 S, 516871 E, 4255884 N

Number of EVA Crew: 3

Participants: Zac Trolley, Tatsunari Tomiyama, Renee Garifi

EVA Commander: Renee Garifi

Rovers used: ATV

1, 2, 3

Duration: 1hours 29minutes

EVA Departure Time: 14:16

EVA Return Time: 15:45

Summary:

The objective of this EVA was to collect meteorite samples at the toothy ridge. We prepared and entered the airlock in timely manner. However, all crew members have issues breathing while walking. Therefore, we decided to return to avoid a critical incident. The details on our breathing issues for crew members will be reported in the operation report. Please reference the operation issue report as a detailed summary.

Sol Summary Report – February 9th

Crew 188 Sol 12 Summary Report 9FEB2018

Sol 12

Summary Title: One more round then it’s [oxygen] bottles to the ground

Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander

Mission Status: EVAs are complete, it’s the final crunch.

Sol Activity Summary:

As the Olympic flame is lit back on Earth, our physical challenges in the field come to a close. With two final EVAs, the crew is dusting off the suits and preparing them for the next crew who will start a week later. The morning EVA hiked up to the Hab Ridge for some final 360 filming, chlorophyll sampling and scenic views. The afternoon EVA had a few samples along the way to Toothy Ridge to the North. Final working photos were snapped with a few last planned for tomorrow during our campus cleanup.

As with any online event on Earth, our Mars-2-Mars event will have a slight change of plans with the session being recorded on WebEx and shared immediately afterwards by the Austrian Space Forum (Österreichisches Weltraum Forum, OeWF). Q&A can be submitted to reddit and the crews will address them during the hangout or afterwards as an ongoing discussion.

10Feb2018, 17:00 CET = 09:00 MDRS time (Sol 13) = 20:00 Oman time [16:00 GMT]

reddit Q&A: https://www.reddit.com/r/Mars/comments/7s23yz/mdrs_crew_188_ama

Spacesuit Up!

Ryan L. Kobrick, Ph.D.

MDRS Crew 188 Commander

Look Ahead Plan:

A hangout with #SimulateMars in Oman, the AMADEE-18 mission at Kepler Station. Then we will end the MDRS Crew 188 simulation mid-day as we continue to clean all areas. With our continuous Hab-chores cleaning schedule we should be on track for an early Sunday morning departure.

Anomalies in work:

None.

Weather:

Calm winds and hazy skies with warm dropping all day.

Crew Physical Status:

There is a mixed of anxiety to get work completed, excitement to return home to loved ones, and sadness to be separating form our Martian family.

EVA:

15 EVAs were completed during our mission.

Reports to file:

1. Ops Report

2. Sol Summary

3. Journalist Report

4. Green Hab Report

5. EVA #14 and #15 report

6. Daily Photos

7. Health & Safety Officer Report

8. Mission Summary

Support Requested (same as previous three reports):

– Heater for science dome requested. A purchase order request was made.

– FYI for mission support to be aware of for systems to track and safe operations.

– There are potentially two packages at Hollow Mountain that we would love to have picked up.

– We will pick them up post-sim

– Vegetables were requested due to the insufficient supply for six crewmembers. Non-vegetarians have cut their personal rations to help stretch the supplies for the two vegetarians. We have two celiac members who cannot eat the starches provided. These diets were all communicated before our arrival. In order to feed crewmembers over the next 4 days, 2-3 cans of dried vegetables were requested (or fresh) with the missing items from the original food supply. We were informed today that we will not be getting any form of resupply. Susan Holden Martin suggested we reiterate this request to mission support since it remains open.

Greenhab Report – February 9th

Crew 188 Sol 12 Green Hab Report 09FEB2018

Julia DeMarines

09/02/2018

Environmental control: Heating

Ambient with door opening: Shade cloth on

Working hours in Green Hab: 13:00 – 15:00, 19:00 – 20:00

Outside temp at working hour: 11ºC

Inside temp at working hour: 14ºC,

Inside temp H: 32ºC,

Inside temp L: 14ºC,

Inside humidity: 35 %

Inside humidity H: 45%

Inside humidity L: 14%

Hours of supplemental light: 06:00 – 09:59

Changes to the crops: N/A

Daily water usage for crops: 2 gallons midday; 10 gallons evening

Time(s) of watering plants: 13:00 – 15:00 (2 gallons), 19:00 – 20:00 (10 gallons)

Changes to research plants: N/A

Narrative: The plants are looking good! We have harvested a lot of dill, leafy greens, and herbs today.

Future needs and questions: We need tomato cages before repotting as to not disturb them any more than necessary.

Summary of NASA VEGGIE OASYS: Renee monitored her experiment with one seed germinating.

EVA Report – February 8th

EVA #13 EVA Summary Follow-up 08Feb2018

Start Time: 10:20

End Time: 13:40

Purpose of EVA: To reconnaissance of Candor Chasma for future EVA mission planning.

Destination: Candor Chasma

UDM27 Coordinates: 12 S 520217 E, 4251607 N

Participants: Ryan Kobrick, Zac Trolley, Sara Jane Pell

Road(s) and routes per MDRS map:

1. Take the entrance road (Driving)

2. Turn left Cow Dung Road (Driving)

3. Turn right Galileo Road (Driving)

4. Turn right Cactus Road (Driving)

5. Park on Cactus Road and walk to the Candor Chasma

6. Return by the same routes.

Mode of Travel: Rover and Pedestrian

Vehicles you will be using (if applicable): Sprit and Opportunity

Mode of travel: Walking and Driving
Vehicles used: Curiosity and Spirit

Since EVA – 13 , several analogue astronauts (of varying height, neck/head girth, lung capacity and fitness) could only maintain a maximal breathing rate for a maximum of a couple of minutes before exhaustion would cause them to reduce their breathing rate, forcing them to suspend their activity, to mitigate CO2 build-up and low visibility condensation/fog inside the helmet. We found that several crew were experiencing the effects of low-oxygen and high-CO2 build up while performing mild walking, even driving vehicles, in warm conditions with low wind.

I t is our assumption that the ambient wind plays a big role in circulating the air within the helmet, and when there is no wind, the internal fans to not have sufficient flow to perform an air exchange at the same rate as the crew members’ breathing level. The fans in the MDRS Analogue Astronaut suits produce 200 Cubic Feet per Minute – this measurement is based on an unencumbered and unrestricted output.

This afternoon we did a full survey of the suit s and their function, and found a systemic fault in the suits. Four of the six suits were running on only one fan. The fan that was on the switch side of the suit has their blades bent and would not rotate. In some cases the fans could be dislodged, and in others the blades were permanently bent. The LEDS still light up, and (half the air) air still flow s through both tubes, so this design flaw was not found until now.